[clas_members] [Revised Logentry] RC Daily Summary

jpoudel at jlab.org jpoudel at jlab.org
Wed Jul 20 10:00:02 EDT 2022

Logentry Text:
----- Tuesday Day Shift (07/19/2022) -----
Asked MCC to stop beam to the hall at 9:00 am for target change NH3 -> ND3
Target change completed around 11:45 am and started polarizing https://logbooks.jlab.org/entry/4017477
Asked MCC to put hall in beam permit. Got beam around 12:45 pm and started production run on ND3
Changed the raster amplitude Vpp from 2.1 V -> 2.2 V
Day shift summary by Stephen Bueltmann: https://logbooks.jlab.org/entry/4017365

----- Tuesday Swing Shift (07/19/2022)-----
Continued production run with ND3 target @4nA until 19:00 pm
No beam after 19:00 pm due to various issues (Storm + injector + ARC3 box)
Swing Shift Summary by Silvia Niccolai: https://logbooks.jlab.org/entry/4017566

----- Wednesday Owl Shift (07/20/2022) -----
No beam at all: https://logbooks.jlab.org/entry/4017706
Shift expert called to inform no beam tonight around 1:15am. Called PDL (Eugene) and target expert (Chris) to ramp down the magnets (solenoid+torus)
Owl Shift crew left CH around 2:15 am (shift cancelled) and Eugene informed Day shift crews to be on standby
Maintenance day started at 6:30 am
Hall in restricted access
Owl Shift Summary by Nicholas Zachariou: https://logbooks.jlab.org/entry/4017695

Plans ahead:
Target survey
iochlb shielding

RC meeting will be at 2:15 pm today. Meeting link: https://jlab-org.zoomgov.com/j/1608674253?pwd=SFZyYmZLWXJieGtpV2dEdVRiUEJnUT09&from=addon#success


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4017801
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220720/62051c9e/attachment.html>

More information about the clas_members mailing list